# Asymptotically conformal CFL quark matter within a nonlocal chiral quark model

### v1

@@ II Nonlocal NJL model for CFLL quark matter
where the subscript index ”$a$” labels quarks ($a=+$) and antiquarks ($a=-$), $\epsilon_{\bf k}^{\pm}=\epsilon_{\bf k}\mp\mu\mp\omega g_{\bf k}$ is expressed through $\epsilon_{\bf k}=\sqrt{{\bf k}^{2}+M_{\bf k}^{2}}$ and $\Delta_{j\bf k}=\zeta_{j}|\Delta|g_{\bf k}$ with $\zeta_{\rm sing}=2$ and $\zeta_{\rm oct}=1$. It is worth mentioning that both the thermodynamic potential (13) and the single particle energies (17) include modulus of $\Delta$. Therefore, for shortening the notations below we suppress the modulus sign when referring to that variable.
It follows from Eq. (17) that the Fermi momentum of quarks $k_{F}$ is defined by the condition $\epsilon_{\bf k}^{+}=0$. Note, for definiteness we consider positive values of the chemical potential. It is also clear that at this momentum the single particle energy of quarks experiences a discontinuous jump. Its amplitude $2\Delta_{j\bf k}|_{|{\bf k}|=k_{F}}$ is twice the gap of the energy spectrum of the corresponding quark state. Since it is defined by $\Delta$ generated by the diquark pairing, in what follows the later is referred to as the diquark pairing gap.
@@
Here $d_{j}^{\pm}=\pm d_{j}$ is introduced for shortening the notations. Unlike the local formulations of chiral quark models (see e.g. Ref. Ivanytskyi and Blaschke (2022a, b)), the previous equations do not include terms with the Dirac delta-function $\delta(\epsilon^{a}_{\bf k})$ arising from the derivative of the sign-function in the dispersion relation (17). In the case of quarks it is provided by $01/2-f^{+}_{j\bf k}=$ 0 at the quark Fermi momentum, while for antiquarks the delta-function vanishes itself due to $\epsilon^{-}_{\bf k}>0$. It is worth mentioning that due to the presence of the formafactor under the momentum integrals Eqs. (23) - (25) do not require any additional regularization of the zero point terms. Their convergence is automatically provided if at large momenta $g_{\bf k}$ decays faster than ${\bf k}^{-2}$. It is important to note, this feature also provides the renormalizability of the model since no additional ultraviolet momentum cutoff is needed in this case.
The solutions of Eqs. (23) - (25) can be used in order to obtain the regularized thermodynamic potential, which gives a direct access to the EoS of the CFLL quark matter. The pressure defined as $p=\Omega^{\rm reg}_{0}-{\Omega}^{\rm reg}$ vanishes in the vacuum ($T=\mu=0$). Below all the vacuum quantities are labelled with the subscript index ”00”. The baryon number, entropy and energy densities can be found using the thermodynamic identities $n_{B}=\partial p/\partial\mu_{B}$, $s=\partial p/\partial T$ and $\varepsilon=\mu_{B}n_{B}+Ts-p$. Here $\mu_{B}=3\mu$ is the baryochemical potential. Below we also analyze the speed of sound defined as the derivative $c_{S}^{2}=dp/d\varepsilon$ found at a constant value of the entropy per baryon $s/n_{B}$. In the zero temperature regime $s=0$ and this last condition is fulfilled automatically leading to $c_{S}^{2}=\partial\ln\mu_{B}/\partial\ln n_{B}$. We will also consider the dimensionless interaction measure $\delta=1/3-p/\varepsilon$, which is nothing but trace of the energy momentum tensor normalized by $3\varepsilon$. It is widely used for discussion of the conformal behavior of quark matter.
@@
$-T\ln\left(1-f_{j\bf k}^{a}\right)=\epsilon_{j\bf k}^{a}f_{j\bf k}^{a}=\epsilon_{j\bf k}^{a}\theta\left(-\epsilon_{j\bf k}^{a}\right).$ (26)
We also use the Gaussian formfactor
@@
where constant $\Lambda$ defines the momentum scale at which the effective current-current interactions are important. This choice assures ultraviolet convergence of the present model and its qualitative agreement with perturbative QCD results on reaching the conformal limit of quark matter. Indeed, at high densities most of the single quark states have momenta well above $\Lambda$. This diminishes the interaction effects and drives the system toward the conformal limit discussed in Section III.
Given the simplification related to the mass degeneration of the quark flavors, we aim at finding reasonable values of the model parameters rather than on their ultimate determination. Thus, the current quark mass is fixed as the mean value of the current masses of $u$- and $d$-quarks from the particle data group booklet Workman et al. (2022), i.e. $m=(m_{u}+m_{d})/2=3.5$ MeV. Furthermore, $\Lambda$ and $G_{S}$ are fixed according to the vacuum values of the effective quark mass at vanishing momentum $M_{0,{\bf k}=0}$ and chiral condensate per flavor $\langle\overline{f}f\rangle_{0}$. The later can be found from the regularized thermodynamic potential as
@@
Depending on the current quark mass and gauge adopted the lQCD reports the vacuum effective mass at zero momentum to lay in the interval 300-500 MeV Skullerud and Williams (2001); Skullerud et al. (2001); Parappilly et al. (2006); Burgio et al. (2012). The QCD sum rules at the renormalization scale of 1 GeV yield $-(229\pm 33~{}\rm MeV)^{3}$ Dosch and Narison (1998) and $-(242\pm 15~{}\rm MeV)^{3}$ Jamin (2002) for the vacuum chiral condensate per flavor. The Gell-Mann-Oakes-Renner relation Gell-Mann et al. (1968) with the pion mass and decay constant from the particle data group booklet Workman et al. (2022) and the current quark mass mentioned above predicts $-(280\pm 12~{}\rm MeV)^{3}$. Therefore, in this work we use the compromised values $M_{0,{\bf k}=0}=400$ MeV and $\langle\overline{f}f\rangle_{0}=-(250~{}\rm MeV)^{3}$. Fitting them yields $\Lambda=564~{}\rm MeV$ and $G_{S}=9.62~{}{\rm GeV}^{-2}$.
We also require the vacuum state to be normal but not color-superconducting. The corresponding condition was found in Ref. Blaschke et al. (2023a). It reads
@@
where the scalar field in the vacuum $\sigma_{0}=M_{0,{\bf k}=0}-m$ satisfies Eq. (23) at $T=\mu=\omega=\Delta=0$. This yields $\eta_{D}<0.765$. In what follows $\eta_{D}$ in this interval and $\eta_{V}$ are considered as the parameters controlling the high density behavior of the model.
@@ III Nonlocal NJL model at high densities
At high densities color superconductivity is the ground state of QCD, while the corresponding pairing gap vanishes only asymptotically Schäfer and Wilczek (1999). At this regime QCD restores the conformal behavior due to the asymptotic freedom and negligebly small quark masses. Before going further, we would like to consider the high density behavior of the CFLL quark matter in order to demonstrate that it respects the mentioned features of QCD. For the sake of generality the following analysis is performed for an arbitrary formfactor. The Gaussian case is considered after the general expressions for the relevant thermodynamic quantities are obtained.
First, one has to prove that once a color-superconducting state is formed in the CFLL quark matter, it preserves itself at arbitrary high densities, which is equivalent to the absence of a high density second order phase transition within the present model. The condition of such a phase transition is $\partial^{2}\Omega^{\rm reg}/\partial\Delta^{2}=0$ at vanishing pairing gap. It can be obtained by dividing Eq. (25) by the gap and later setting $\Delta=0$. The quark contribution to the right hand side part of this condition diverges due to the factor $(\epsilon_{\bf k}-\mu)^{-1}$ under the momentum integral. This divergence is naturally regularised in the presence of a non-zero pairing gap. Consequently, this right hand side part diverges at vanishing pairing gap and, thus, is not equal to the unit in the left hand side part. This allows us to conclude that the condition of a second order phase transition at high densities can not be fulfilled, which implies the pairing gap is non-zero at that regime. This proves that, alike QCD, the color superconductivity is the ground state of cold and dense quark matter. At the same time, at finite or even diverging pairing gap the right hand side part of Eq. (25) vanishes at high chemical potentials due to the factor $\epsilon^{a}_{j\bf k}$ under the momentum integral and, thus, is not equal to $\Delta$ in its left hand side part. Thus, the pairing gap of the CFLL quark matter asymptotically vanishes in the high density regime.
This allows us to approximate $\epsilon_{\bf k}^{a}/\epsilon_{j\bf k}^{a}\simeq 1$ in Eqs. (23) - (24). Furthermore, the single particle distribution function of quarks can be decomposed as $f_{j\bf k}^{+}=1-\theta(|{\bf k}|-k_{F})$, where the first and the second terms represent the dominant contribution and the leading order corrections to the momentum integrals. Using this decomposition and performing the summation over the quark-antiquark states we obtain the leading terms of the scalar and vector fields
@@
where $\tilde{d}=d_{\rm sing}\zeta_{\rm sing}^{2}+d_{\rm oct}\zeta_{\rm oct}^{2}$. It is interesting to compare this results to the one gluon exchange perturbative calculations Schäfer and Wilczek (1999). They suggest the quark pairing gap $\Delta_{\rm QCD}\propto\mu g_{\rm QCD}^{-5}\exp(-3\pi^{2}/\sqrt{2}g_{\rm QCD})$ with $g_{\rm QCD}$ being the QCD coupling. Equating $\Delta_{\infty}$ and $\Delta_{\rm QCD}$ brings the high density behavior of the formfactor $g_{\mu}\propto\sqrt{g_{\rm QCD}}/\mu$. Running of the QCD coupling can be estimated according to the one-loop $\beta$-function as $g_{\rm QCD}\propto\ln^{-1/2}(\mu^{2}/\Lambda^{2}_{\rm QCD})$. From this we conclude that agreement with the perturbative QCD calculations requires the formfactor to behave at high momenta as $|{\bf k}|^{-1}$. Such a behavior does not provide convergence of the momentum integrals in Eqs. (23) - (25) and spoils the renormalizability of the nonlocal NJL model. Requiring such a renormalizability disables determination of the formafactor of the nonlocal NJL model from the perturbative QCD calculations of the quark pairing gap.
In the Gaussian case considered in this paper the momentum intergals in Eqs. (30) and (31) can be performed explicitly. This yields the asymptotics of the vector and scalar fields $\sigma_{\infty}=dG_{S}\Lambda^{2}m\exp(-\mu^{2}/\Lambda^{2})/2$ and $\omega_{\infty}=-dG_{V}\Lambda^{3}/4\pi^{3/2}$. The pairing gap decays double exponentially with $\mu^{2}$ in this case.
Finding the asymptotics of the speed of sound and dimensionless interaction measure of cold CFLL matter requires accounting for the fact that quark Fermi momentum is reduced by the vector repulsion and quark mass. Saturation of the vector field to the constant value found above and convergence of the effective quark mass to the current one brings us to the conclusion that at high densities $k_{F}\simeq\mu+\omega_{\infty}g_{\mu}-m^{2}/2\mu$. For the formfactors providing convergence of the momentum integrals in Eqs. (23)-(25) the quark mass correction to $k_{F}$ dominates against the vector repulsion. This, however, is not the case in the chiral limit when $m=0$. Therefore, both the mass and vector repulsion corrections are kept for the sake of generality of the analysis. At high densities $n_{B}\simeq dk_{F}^{3}/18\pi^{2}$. Using this, we arrive at
@@
The asymptotic of the dimensionless interaction measure can be found by noticing that at high densities the pressure to energy density ratio is an indeterminate form $\infty$/$\infty$ / ∞, which can be treated using the L’Hôpital’s rule. This along with the thermodynamic identities mentioned above yields $\delta=1/3-c_{S}^{2}$. Note, this model independent relation holds only at asymptotically high densities but applies even for the systems, which do not converge to the conformal limit.
Thus, the CFLL quark matter asymptotically approaches the conformal limit $c_{S}^{2}=1/3$, $\delta=0$ due to vanishing of the formfactor and negligibly small quark mass. Such an asymptotically conformal behavior was reported within the confining density functional approach for color-superconducting quark matter with medium dependent couplings Ivanytskyi and Blaschke (2022b); Gärtlein et al. (2023). Within that approach the conformal values of the speed of sound and dimensionless interaction measure are reached from above and below, respectively. At the same time, the perturbative QCD calculations for cold and dense quark matter indicate reaching them from the opposite sides. It follows from Eq. (34) that given a negative value of $\omega_{\infty}$, the conformal limit within the present model is reached in a qualitative agreement with perturbative QCD for any formactor providing convergence of the momentum integrals in Eqs. (23) - (25) regardless the value of $m$. However, the formafactor can not be chosen to simultaneously provide the renormalizability of the present model and reproduce the two-loop perturbative result assuming that deviation of speed of sound from the conformal value scales as $g_{\rm QCD}^{4}$ Blaschke et al. (2023b). This signals about an inherent non-perturbative charcter of the nonlocal NJL model.
It is important to stress that the mentioned qualitative agreement of the considered model with the predictions of perturbative QCD is provided by the nonlocal character of the quark-quark interactions. In order to demonstrate this we want to compare the present model to the version of the NJL one from Ref. Blaschke et al. (2023a). Similarly to this work, that model treats the scalar and diquark interaction channels nonlocally and adopts for them the Gaussian formfactor. At the same time, the vector interaction of that version of the NJL model is local. The corresponding thermodynamic potential and mean-field equations can be obtained by replacing $\omega g_{\bf k}$ by $\omega$ where applicable and setting the formafactor in Eq. (24) equal to unit. This, however, does not change the conclusion about vanishing $\sigma_{\infty}$ and $\Delta_{\infty}$, which holds due to the nonlocal treatment of the corresponding interaction channels. Then, only the contribution of quarks survives in Eq. (24) with $g_{\bf k}=1$. This yields

### v2

@@ II Nonlocal NJL model for CFLL quark matter
where the superscript index ”$a$” labels quarks ($a=+$) and antiquarks ($a=-$), $\epsilon_{\bf k}^{\pm}=\epsilon_{\bf k}\mp\mu\mp\omega g_{\bf k}$ is expressed through $\epsilon_{\bf k}=\sqrt{{\bf k}^{2}+M_{\bf k}^{2}}$ and $\Delta_{j\bf k}=\zeta_{j}|\Delta|g_{\bf k}$ with $\zeta_{\rm sing}=2$ and $\zeta_{\rm oct}=1$. It is worth mentioning that both the thermodynamic potential (13) and the single particle energies (17) include modulus of $\Delta$. Therefore, for shortening the notations below we suppress the modulus sign when referring to that variable.
It follows from Eq. (17) that the Fermi momentum of quarks $k_{F}$ is defined by the condition $\epsilon_{\bf k}^{+}=0$. Note, for definiteness we consider positive values of the chemical potential. It is also clear that at this momentum the single particle energy of quarks experiences a discontinuous jump. Its amplitude $2\Delta_{j\bf k}|_{|{\bf k}|=k_{F}}$ is twice the gap of the energy spectrum of the corresponding quark state. Since it is defined by $\Delta$ generated by the diquark pairing, in what follows the latter is referred to as the diquark pairing gap.
@@
Here $d_{j}^{\pm}=\pm d_{j}$ is introduced for shortening the notations. Unlike the local formulations of chiral quark models (see e.g. Ref. Ivanytskyi and Blaschke (2022a, b)), the previous equations do not include terms with the Dirac delta-function $\delta(\epsilon^{a}_{\bf k})$ arising from the derivative of the sign-function in the dispersion relation (17). In the case of quarks it is provided by $01/2-f^{+}_{j\bf k}=$ 0 at the quark Fermi momentum, while for antiquarks the delta-function vanishes itself due to $\epsilon^{-}_{\bf k}>0$. It is worth mentioning that due to the presence of the formafactor under the momentum integrals Eqs. (23) - (25) do not require any additional regularization of the zero point terms. Their convergence is automatically provided if at large momenta $g_{\bf k}$ decreases with growth of $\bf k$ faster than ${\bf k}^{-2}$. It is important to note, this feature also provides the renormalizability of the model since no additional ultraviolet momentum cutoff is needed in this case.
The solutions of Eqs. (23) - (25) can be used in order to obtain the regularized thermodynamic potential, which gives a direct access to the EoS of the CFLL quark matter. The pressure defined as $p=\Omega^{\rm reg}_{0}-{\Omega}^{\rm reg}$ vanishes in the vacuum ($T=\mu=0$). Below all the vacuum quantities are labeled with the subscript index ”00”. The baryon number, entropy and energy densities can be found using the thermodynamic identities $n_{B}=\partial p/\partial\mu_{B}$, $s=\partial p/\partial T$ and $\varepsilon=\mu_{B}n_{B}+Ts-p$. Here $\mu_{B}=3\mu$ is the baryochemical potential. Below we also analyze the speed of sound defined as the derivative $c_{S}^{2}=dp/d\varepsilon$ evaluated at a constant value of the entropy per baryon $s/n_{B}$. In the zero temperature regime $s=0$ and this last condition is fulfilled automatically leading to $c_{S}^{2}=\partial\ln\mu_{B}/\partial\ln n_{B}$. We will also consider the dimensionless interaction measure $\delta=1/3-p/\varepsilon$, which is nothing but trace of the energy momentum tensor normalized by $3\varepsilon$. This quantity is widely used for discussion of the conformal behavior of quark matter.
@@
$T\ln\left(1-f_{j\bf k}^{a}\right)=\epsilon_{j\bf k}^{a}f_{j\bf k}^{a}=\epsilon_{j\bf k}^{a}\theta\left(-\epsilon_{j\bf k}^{a}\right).$ (26)
In what follows we use the Gaussian formfactor
@@
where constant $\Lambda$ defines the momentum scale below which the effective current-current interactions are important. This choice guarantees ultraviolet convergence of the present model and its qualitative agreement with perturbative QCD results in reaching the conformal limit of quark matter. At high densities most of the single quark states have momenta well above $\Lambda$. This diminishes the interaction effects and drives the system toward the conformal limit discussed in Section III.
The formfactor can be chosen in other functional forms, e.g. in the Lorentzian-type or Woods-Saxon ones Grigorian (2007). However, at non-perturnative regimes typical for the NS interiors specific choice of $g_{\bf k}$ is not too important since $k_{F}<\Lambda$ and for the momentum states below the Fermi sphere all the mentioned functional forms of the formfactor differ by not more than 25%. This makes modelling NS within the framework of the nonlocal NJL model weakly sensitive to the formfactor choice.
Given the simplification related to the mass degeneration of the quark flavors, we aim at finding reasonable values of the model parameters rather than at their ultimate determination. Thus, the current quark mass is fixed as the mean value of the current masses of $u$- and $d$-quarks from the Review of Particle Physics Workman et al. (2022), i.e. $m=(m_{u}+m_{d})/2=3.5$ MeV. Furthermore, $\Lambda$ and $G_{S}$ are fixed according to the vacuum values of the effective quark mass at vanishing momentum $M_{0,{\bf k}=0}$ and chiral condensate per flavor $\langle\overline{f}f\rangle_{0}$. The latter can be found from the regularized thermodynamic potential as
@@
Depending on the current quark mass and gauge adopted in the lQCD simulations, the vacuum effective mass at zero momentum is found to lay in the interval 300-500 MeV Skullerud and Williams (2001); Skullerud et al. (2001); Parappilly et al. (2006); Burgio et al. (2012). The QCD sum rules at the renormalization scale of 1 GeV yield $-(229\pm 33~{}\rm MeV)^{3}$ Dosch and Narison (1998) and $-(242\pm 15~{}\rm MeV)^{3}$ Jamin (2002) for the vacuum chiral condensate per flavor. The Gell-Mann-Oakes-Renner relation Gell-Mann et al. (1968) with the pion mass and decay constant from the Review of Particle Physics Workman et al. (2022) and the current quark mass mentioned above predicts $-(280\pm 12~{}\rm MeV)^{3}$. In this work we use the compromise values $M_{0,{\bf k}=0}=400$ MeV and $\langle\overline{f}f\rangle_{0}=-(250~{}\rm MeV)^{3}$. Fitting them yields $\Lambda=564~{}\rm MeV$ and $G_{S}=9.62~{}{\rm GeV}^{-2}$.
Figure 1: Scalar field $\sigma$ (upper left panel), vector field $\omega$ (upper right panel), pairing gap $\Delta$ (lower left panel) and critical value of the strange quark mass $M_{s}^{c}$ as functions of baryon chemical potential $\mu_{B}$ obtained for the vector and diquark couplings presented in Table 1. The thin curves on the lower right panel correspond to the effective mass of the zero momentum strange quarks $M_{s,{\bf k}=0}$ calculated at $m_{s}=95$ MeV. The empty white circles indicate the first order phase transition in the hybrid quark-hadron EoS (see Section IV for details). The filled colored circles represent the maximum chemical potentials in the centers of the heaviest NSs with the corresponding EoSs.
We also require the vacuum state to be normal, i.e. not color-superconducting. The corresponding condition was found in Ref. Blaschke et al. (2023a). It reads
@@
where the scalar field in the vacuum $\sigma_{0}=M_{0,{\bf k}=0}-m$ satisfies Eq. (23) at $T=\mu=\omega=\Delta=0$. This yields $\eta_{D}<0.765$. In what follows $\eta_{D}$ below the mentioned value and $\eta_{V}$ are considered as the parameters controlling the high density behavior of the model.
$\eta_{V}$ $\eta_{D}$ $\rm\varepsilon_{\rm max}$ $\rm M_{\rm max}$ $\rm R_{1.4}$ $\Lambda_{1.4}$
$\rm[MeV/fm^{3}]$ $\rm M_{\rm max}~{}[M_{\odot}]$ $\rm R_{1.4}$ [km] $\Lambda_{1.4}$
0.55 0.27 1281 1.94 12.23 280
0.70 0.34 1203 2.01 12.39 321
0.85 0.38 1164 2.09 12.58 393
1.00 0.41 1086 2.15 12.80 424
1.21 0.44 1038 2.24 13.15 569
Table 1: Vector $\eta_{V}$ and diquark $\eta_{D}$ couplings producing the solutions of the mean-field equations shown of Fig. 1, EoSs shown in Fig. 2, maximal energy densities $\varepsilon_{\rm max}$ reached in the centers of the heaviest NSs, as well as the corresponding maximum masses $\rm M_{\rm max}$, radii $\rm R_{1.4}$ and tidal deformabilities $\Lambda_{1.4}$ of a canonical NS extracted from Figs. 4 and 5.
Fig. 1 shows solutions of the coupled mean-field equations (23) - (25) solved for the values of the model parameters specified above and for the vector and diquark couplings given in Table 1. At small chemical potentials the vector field and pairing gap vanish, while the scalar field has its vacuum value. This corresponds to the chirally broken normal phase of cold QCD matter, which is the ground state at small $\mu_{B}$. Within the MFA baryon density and pressure of this phase vanish since quarks have large effective masses and can not be excited. Going beyond the MFA introduces to the consideration nucleons and hyperons, which generate finite baryon densities and pressures even in the absence of quarks. In Section IV this is modeled by constructing a hybrid quark-hadron EoS. At high chemical potentials scalar field deviates from its vacuum value and the pairing gap is finite. This signals about the partially chirally restored color-superconducting phase of cold QCD matter, the CFLL phase. This phase is the ground state of QCD at high densities and is constituted by the Cooper pair type diquark correlations of quarks with small effective masses. The vector field, baryon density and pressure attain finite values in this phase. The transition between the chirally broken normal and partially chirally restored color-superconducting phases of cold QCD matter occurs when their pressures coincide. As is seen from Fig. 1, depending on the values of the vector and diquark couplings the order parameters of chiral symmetry ($\sigma$) and color superconductivity ($\Delta$) change continuously or discontinuously at this transition. At large $\eta_{V}$ and $\eta_{D}$ this change is continuous but exhibits kinks in the behavior of $\sigma$ and $\Delta$, which signals about a second order phase transition. At small values of the vector and diquark couplings the mentioned order parameters change discontinuously and the transition is of the first order.
While the present model neglects the mass splitting between different quark flavors, the quark pairing patterns can be modified if the current mass of strange quarks has a physical value $m_{s}\gg m$. More specifically, at low densities heavy strange quarks are not gapped if their effective mass $M_{s,{\bf k}}=m_{s}+\sigma g_{\bf k}$ exceeds a certain critical value $M_{s}^{c}$ Schäfer and Wilczek (1999b); Alford et al. (1999b); Abuki (2003). Here the current quark mass has its physical value $m_{s}=95$ MeV Workman et al. (2022) and $\sigma$ is calculated at $m_{s}=m$. Under the mentioned condition strange quarks are unlocked from the CFL phase of quark matter, which corresponds to the 2SC quark matter Buballa (2005). The critical value of the effective mass of strange quarks depends on the strength of the diquark pairing interaction and density of quark matter. Based on a simple kinematic consideration it reads Schäfer and Wilczek (1999b); Alford et al. (1999b); Abuki (2003)
$M_{s}^{c}=2\sqrt{k_{F}\Delta},$ (30)
where the quark Fermi momentum and pairing gap are calculated at $m_{s}=m$. Fig. 1 shows the behavior of $M_{s}^{c}$ and the effective mass of the zero momentum strange quarks $M_{s,{\bf k}=0}$. Note, at finite momenta $M_{s,{\bf k}}<M_{s,{\bf k}=0}$. It is seen from Fig. 1 that for the considered values of the vector and diquark couplings and at small values of the baryonic chemical potentials below about 1000 GeV $M_{s,{\bf k=}0}>M_{s}^{c}$. In this range of $\mu_{B}$ the simplification related to mass degeneration of quarks is inapplicable, since it overlooks the transition between the CFL and 2SC phases of quark matter Blaschke et al. (2005, 2009, 2010). At the same time, $M_{s}^{c}$ always exceeds $M_{s,{\bf k}=0}$ above the first order phase transition in the hybrid quark-hadron EoS described in Section IV. This means that accounting for the physical current mass of strange quarks within the present model does not change the CFL pattern of quark pairing and does not bring it to the 2SC one if $\mu_{B}$ exceeds about 1000 MeV. Furthermore, for the vector and diquark couplings providing a good agreement with the observational data on NS ($\eta_{V}\geq 0.7$, see Section IV for details) and $\mu_{B}>1200$ MeV $M_{s}^{c}$ exceeds $M_{s,{\bf k}=0}$ at least by a factor about two. This allows us to conclude that in the mentioned range of the baryonic chemical potentials accounting for a physical value of $m_{s}$ produces a small correction to the EoS of quark matter that can be accounted for perturbatively in the spirit of the phenomenological parametrization by Alford, Braby, Paris and Reddy Alford et al. (2005).
@@ III Nonlocal NJL model at high densities
At high densities color superconductivity is the ground state of QCD, while the corresponding pairing gap vanishes only asymptotically Schäfer and Wilczek (1999a). At this regime QCD restores the conformal behavior due to the asymptotic freedom and negligebly small quark masses. Before going further, we consider the high density behavior of the CFLL quark matter in order to demonstrate that it respects the mentioned features of QCD. For the sake of generality the following analysis is performed for an arbitrary formfactor. The Gaussian case is considered after the general expressions for the relevant thermodynamic quantities are obtained.
First, one has to prove that once a color-superconducting state is formed in the CFLL quark matter, it preserves itself at arbitrary high densities, which is equivalent to the absence of a high density second order phase transition within the present model. The condition of such a phase transition is $\partial^{2}\Omega^{\rm reg}/\partial\Delta^{2}=0$ at vanishing pairing gap. Dividing Eq. (25) by the gap and later setting $\Delta=0$ yields this condition in the explicit form. The quark contribution to its right hand side part diverges due to the factor $(\epsilon_{\bf k}-\mu)^{-1}$ under the momentum integral. This divergence is naturally regularised in the presence of a non-zero pairing gap. Consequently, this right hand side part diverges at vanishing pairing gap and, thus, is not equal to the unity in the left hand side part. This allows us to conclude that the condition of a second order phase transition at high densities can not be fulfilled and the pairing gap is non-zero. This proves that, just like in QCD, the color superconductivity is the ground state of cold and dense quark matter. At the same time, at high densities Eq. (25) can not be satisfied by finite or even diverging pairing gap. In this case the right hand side part of Eq. (25) vanishes at high chemical potentials due to the factor $\epsilon^{a}_{j\bf k}$ under the momentum integral and, thus, is not equal to $\Delta$ in its left hand side part. In other words, fulfilling Eq. (25) requires $\Delta\rightarrow 0$ at $\mu\rightarrow\infty$. Thus, the pairing gap of the CFLL quark matter asymptotically vanishes in the high density regime.
This allows us to approximate $\epsilon_{\bf k}^{a}/\epsilon_{j\bf k}^{a}\simeq 1$ in Eqs. (23) - (24), while the subdominant order $\mathcal{O}(\Delta^{2})$ correction can be neglected. Furthermore, the single particle distribution function of quarks can be decomposed as $f_{j\bf k}^{+}=1-\theta(|{\bf k}|-k_{F})$, where the first and the second terms represent the dominant contribution and the leading order corrections to the momentum integrals. Using this decomposition and performing the summation over the quark-antiquark states we obtain the leading terms of the scalar and vector fields
@@
where $\tilde{d}=d_{\rm sing}\zeta_{\rm sing}^{2}+d_{\rm oct}\zeta_{\rm oct}^{2}$. It is interesting to compare this results to the one gluon exchange perturbative calculations Schäfer and Wilczek (1999a). They suggest the quark pairing gap $\Delta_{\rm QCD}\propto\mu g_{\rm QCD}^{-5}\exp(-3\pi^{2}/\sqrt{2}g_{\rm QCD})$ with $g_{\rm QCD}$ being the QCD coupling. Equating $\Delta_{\infty}$ and $\Delta_{\rm QCD}$ brings the high density behavior of the formfactor $g_{\mu}\propto\sqrt{g_{\rm QCD}}/\mu$. Running of the QCD coupling can be estimated according to the one-loop $\beta$-function as $g_{\rm QCD}\propto\ln^{-1/2}(\mu^{2}/\Lambda^{2}_{\rm QCD})$. From this we conclude that agreement with the perturbative QCD calculations requires the formfactor to behave at high momenta as $|{\bf k}|^{-1}$. This behavior of the formfactor leads to divergent momentum integrals in Eqs. (23) - (25). At $g_{\bf k}\propto|{\bf k}|^{-1}$ their regularization requires an ultraviolet momentum cutoff, which spoils the renormalizability of the nonlocal NJL model. Requiring the renormalizability disables using the formafactor extracted from the perturbative QCD calculations of the quark pairing gap.
In the Gaussian case considered in this paper the momentum integrals in Eqs. (31) and (32) can be performed explicitly. This yields the asymptotes of the vector and scalar fields $\sigma_{\infty}=dG_{S}\Lambda^{2}m\exp(-\mu^{2}/\Lambda^{2})/2$ and $\omega_{\infty}=-dG_{V}\Lambda^{3}/4\pi^{3/2}$. For the considered formfactor the pairing gap decreases double exponentially at growing $\mu^{2}$.
Finding the asymptote of the speed of sound and dimensionless interaction measure of cold CFLL matter requires accounting for the fact that quark Fermi momentum is reduced by the vector repulsion and quark mass. Saturation of the vector field to the constant value found above and convergence of the effective quark mass to the current mass brings us to the conclusion that at high densities $k_{F}\simeq\mu+\omega_{\infty}g_{\mu}-m^{2}/2\mu$. For the formfactors providing convergence of the momentum integrals in Eqs. (23)-(25) the quark mass correction to $k_{F}$ dominates against the vector repulsion. This, however, is not the case in the chiral limit when $m=0$. Therefore, both the mass and vector repulsion corrections are kept for the sake of generality of the analysis. At high densities $n_{B}\simeq dk_{F}^{3}/18\pi^{2}$. Using this, we arrive at
@@
The asymptote of the dimensionless interaction measure can be found by noticing that at high densities the pressure to energy density ratio is of an indeterminate form $\infty$/$\infty$ / ∞, which can be treated using the L’Hôpital’s rule. This along with the thermodynamic identities mentioned above yields $\delta=1/3-c_{S}^{2}$. Note that this model independent relation holds only at asymptotically high densities but applies even for the systems, which do not converge to the conformal limit.
Thus, the CFLL quark matter asymptotically approaches the conformal limit $c_{S}^{2}=1/3$, $\delta=0$ due to vanishing of the formfactor and negligibly small quark mass. Such an asymptotically conformal behavior was reported within the confining density functional approach for color-superconducting quark matter with medium dependent couplings Ivanytskyi and Blaschke (2022b); Gärtlein et al. (2023). Within that approach the conformal values of the speed of sound and dimensionless interaction measure are reached from above and below, respectively. At the same time, the perturbative QCD calculations for cold and dense quark matter indicate reaching them from the opposite sides. It follows from Eq. (35) that given a negative value of $\omega_{\infty}$, the conformal limit within the present model is reached in a qualitative agreement with perturbative QCD for any formfactor providing convergence of the momentum integrals in Eqs. (23) - (25) regardless the value of $m$. However, the formfactor can not be chosen to simultaneously provide the renormalizability of the present model and reproduce the two-loop perturbative result assuming that deviation of speed of sound from the conformal value scales as $g_{\rm QCD}^{4}$ Blaschke et al. (2023b). This signals about an inherent non-perturbative character of the nonlocal NJL model.
It is important to stress that the mentioned qualitative agreement of the considered model with the predictions of perturbative QCD is provided by the nonlocal character of the quark-quark interactions. In order to demonstrate this we want to compare the present mode to the version of the NJL model from Ref. Blaschke et al. (2023a). Similarly to this work, that model treats the scalar and diquark interaction channels nonlocally and adopts for them the Gaussian formfactor. At the same time, the vector interaction of that version of the NJL model is local. The corresponding thermodynamic potential and mean-field equations can be obtained by replacing $\omega g_{\bf k}$ by $\omega$ where applicable and setting the formafactor in Eq. (24) equal to unit. This, however, does not change the conclusion about vanishing $\sigma_{\infty}$ and $\Delta_{\infty}$, which holds due to the nonlocal treatment of the corresponding interaction channels. Then, only the contribution of quarks survives in Eq. (24) with $g_{\bf k}=1$. This yields